# Graded lithium-ion battery pouch cells to homogenise current distributions and mitigate lithium plating

### v1

@@ Abstract
Spatial distributions in current, temperature, state-of-charge and degradation across the plane of large format lithium-ion battery pouch cells can significantly impact their performance, especially at high C-rates. In this paper, a method to smooth out these spatial distributions by grading the electrode microstructure in-the-plane is proposed. A mathematical model of a large format pouch cell is developed and validated against both temperature and voltage experimental data. An analytic solution for the optimal graded electrode that achieves a uniform current distribution across the pouch cell is then derived. The model predicts that the graded electrodes could significantly reduce the likelihood of lithium plating in large format pouch cells, with grading increasing the C-rate at which plating occurs from 2.75C to 5.5C. These results indicate the potential of designing spatially varying electrode architectures to homogenise the response of large format pouch cells and improve their high rate performance.
@@ 1 Introduction
Lithium-ion battery electrodes are complex non-linear dynamical systems reacting non-uniformly in space and time, especially when subject to high C-rate currents. Phenomena such as travelling wave fronts in charging LiFePO4 (LFP) cathodes 1; 2 and localised lithium-plating in graphite anodes 3 are just some examples of the diverse range of non-linear behaviours that have been observed in battery electrodes, and it is these localised behaviours that often govern cell performance. To protect against these spatially localised electrochemical behaviours, the use of graded electrode structures (as in those with controlled spatial variations in electrode microstructure) has been proposed with the idea being that introducing heterogeneity into the electrode microstructure can reduce the heterogeneity in the electrochemical response.
@@
Recently, experimental data have provided support for the use of graded electrodes in practical cells 4. A common theme is the potential of graded electrodes to increase both the capacity at high C-rates 5; 6 and the cycle life of thick electrodes 7; 8, with other benefits including a reduction in electrode tortuosity 9. Data for both standard electrodes, such as LFP cathodes 7; 8, and non-standard electrodes, such as anodes arranged in a layered structure composed of high power Li4Ti5O12 and high capacity SnO2 10, have shown similar benefits. However, care has to be taken when designing graded electrodes, as structuring the electrode incorrectly can deteriorate performance 11; 8. To prevent this, mathematical models have been developed to predict the electrochemical response of graded electrodes so as to quickly and cost-effectively optimise their designs. Several different models have now been developed covering a range of length scales (including micro-structural12; 13 and Doyle-Fuller-Newman-type models14; 2) and performance metrics (such as fast charging capacitance 2 and electrical impedance 15). These results highlight the growing maturity of battery modelling and parameterisation, which, in turn, is making it increasingly straightforward to embed modelling pro-actively in the battery design process 16. An example of this battery co-design process are the trapezoidal electrodes7 that were developed using insights 15 on how the distribution of carbon and binder affects the conductivity of the LiFePO4 electrodes.
@@
The focus of most existing studies on graded electrodes for lithium-ion batteries has been on controlling the through-thickness local electrode microstructure 12; 17, including bilayer cathodes2 composed of discrete layers of both     Li[Ni0.6Co0.2Mn0.2]O2 (NMC622) and LFP. Focusing on the through-thickness electrode structure (with experiments typically only involving coin cells 7) is justified by the fact that it is charge mobility through the electrode thickness between the anode and cathode that governs coin cell performance. However, in recent years, there has been a general trend in commercial cells towards larger cell formats (demonstrated, for example, by the introduction of the larger 4680 cylindrical cells 18). The idea being that larger cell formats increase the specific energy density, as the ratio of active material to superfluous materials such as casings and control circuits is increased. Large format cells have their own problems compared to coin cells; in particular, they experience new forms of spatial heterogeneities, which emerge at the macro-scale (as in across their width and length). In particular, thermal hot spots 19; 20 and spatially localised distributions of electrode degradation 21; 22; 23, lithium plating 24, state-of-charge 25, and currents in-the-plane of large-format pouch cells 26; 27 have been identified as impacting performance, especially in high rate applications.
@@
The significance of in-the-plane spatial heterogeneities on the performance of large format pouch cells is becoming increasingly apparent. In turn, this has driven efforts to understand how these heterogeneities develop using mathematical models. Through these models, predictions for chemical reaction 28 and current distributions along the plane of the current collectors 29 (and how these distributions evolve with changes in the cell’s design, e.g. its current collector thickness 30 and tapering 31) have been developed, with analytical solutions even being proposed 32. The push to understand the planar dynamics of pouch cells has meant that the number of spatial dimensions in these models has increased, with 2D 33 and 3D 34; 35; 36; 37; 20; 38 effects often included, as opposed to the through-thickness models of, for example, the DFN model and single particle models 39; 40. One of the main drawbacks of large format battery models is their complexity, a consequence of their multi-dimensional spatial domains. This model complexity leads to several challenges including long simulation times and difficulties in model parameterisation, as discussed in the electro-thermal modelling studies 1920. These challenges demonstrate a need for easily parameterised and computationally efficient pouch cell models able to capture the spatial distributions seen in experimental data. This problem is considered with the model introduced in Section 2.
@@
(a) Representation of electrode/separator sandwich of the pouch cell. The cathode resistance is assumed to be graded. A left handed co-ordinate system is used.
@@
Figure 1: Schematic of the 2D model for a lithium-ion battery pouch cell.
The formation of spatial heterogeneities across pouch cells indicates that it may be possible to grade the local electrode microstructure in-the-plane to mitigate their effect, similar to what has been achieved with through-thickness grading 41; 17. This graded electrode problem is illustrated in Figure 1, with Figure 1(a) indicating the planar and through-thickness directions of the pouch cell, Figure 1(b) indicates the simplification that allows the model to be reduced to a 2D problem, and Figure 1(c) indicates the 2D model domain.
@@
    An analytic solution for the distribution of carbon black in the electrode plane that achieves a flat current distribution across the pouch cell is derived.
@@
    Compared to uniform electrodes, it is predicted that graded electrodes can increase the charging rate when lithium plating occurs from 2.75C to 5.2C.
It is proposed that a primary advantage of these graded pouch cells is to mitigate the impact of lithium plating (see Section 4), as the flatter current distributions reduce the likelihood of negative anode overpotentials. The results highlight the importance of in-the-plane effects when characterising lithium plating in large format cells. By contrast, it is shown that the high thermal/electronic conductivity of the current collectors means that only minimal benefits of grading are observed when considering the thermal and voltage response; Section 3.4 shows that the graded cells only delivered a 1.2% increase in stored capacitance for 4C charging compared to the uniform electrodes, a conclusion in agreement with Hosseinzadeh, et al. 201841. These validated modelling results indicate the potential of grading to improve the cycle-life and high-rate performance of large format pouch cells. More broadly, they indicate how improvements in battery design and manufacturing could reduce the performance gap observed between coin cells constructed in laboratories and the large format cells used in commercial high-rate applications 42.
@@ 2 Mathematical model
In this section, a mathematical model of a large format pouch cell is introduced. Specific focus is given to a comparison between the model and experimental data20, with the presented model being simpler in form (specifically, here, circuit dynamics are used to capture the through-thickness electrical response, whereas a simple electrochemical model is used in Lin, et al, 202220) whilst still being able to accurately capture the in-plane spatial distributions seen in the experimental data (see Section 2.5). Keeping the model formulation relatively simple while retaining the important physics-based mechanisms allows analytic solutions to the optimal graded electrode design problem (explored in Section 3) to be obtained, which is the focus of this work.
@@ 1. A1:
    The current collectors’ thicknesses are much thinner than their width and height.
@@ 3. A3:
    The multi-layer pouch cell can be modelled as an effective single thick layer19; 20.
A1 is justified because large format pouch cells are considered (with typical current collector thicknesses being $\approx 10\mu$m and their planar dimensions being $\approx$ 15cm $\times$ 20cm). Defining $i_{x}(y,z,t)$ as the charging current through the electrode, and $I_{x,k}(x,y,z,t)$ as the current flowing in the $x$-direction through current collector $k$, then, under this thin plate assumption, the current gradient in the $x$-direction through the current collectors can be approximated by
@@
following the ideas of Taheri, et al. 201432. A2 is a simplification of the electrodes’ electrical dynamics, but, even so, the resulting model is able to capture the spatial distributions seen in the experimental data (see Section 2.5) while being simple enough to generate an analytic solution to the optimal electrode design problem (Section 3.1) and parameterise (illustrated by the relative lack of parameters in Table A2). A3 is justified under the assumption that each layer has a similar composition and so react similarly.
@@ 2.2 Model equations
The following equations characterise the proposed mathematical model for large-format pouch cells, with the models domain described in Figure 1(c). Regarding the notation, each model variable in Eqn. (3) is defined in the ($y,z$)-plane of the pouch cell, as illustrated in Figure 1(c), and may also evolve in time $t$. This $(y,z,t)$ dependency of the variables is dropped in the model equations (3) to ease readability. $V(y,z,t)$ is the voltage (defined as the potential of the current collector at the cathode minus that of the anode) at each point in the plane, $T(y,z,t)$ is the cell temperature, $i_{x}(y,z,t)$ is the local through-thickness current, $R(y,z)$ is the local resistance of the two electrodes and the separator that is constant with time, while SoC$(y,z,t)$ is the local state of charge in the plane. Table A1 summarises the model’s variables, while Table A2 gives its parameters. For the graded cell, the only parameter varying in $(y,z)$ is $R(y,z)$, with the rest being constants. Using the notation $\Delta f=\frac{\partial^{2}f}{\partial y^{2}}+\frac{\partial^{2}f}{\partial z^{2}}$ and defining
@@
$i_{x}$ $=\frac{V-U(\text{SoC,T})-v_{1}-v_{2}}{R},$ (3a)
$\Delta V$ $=-\frac{i_{x}}{\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})},\ \ \text{on}\ \ \Omega,\ t>0,$ (3b)
$\rho\frac{\partial T}{\partial t}$ $=\lambda_{\scriptscriptstyle\text{eff}}\Delta T+R\frac{{i_{x}}^{2}\ }{L_{x}}-h_{\scriptscriptstyle\text{faces}}\left(T-T_{\scriptscriptstyle\text{ref}}\right),\ \ \text{on}\ \ \Omega,\ t>0,$ (3c)
$\frac{\partial\text{SoC}}{\partial t}$ $=\frac{i_{x}}{C_{\scriptscriptstyle\text{cell}}},\ \ \text{on}\ \ \Omega,\ t>0,$ (3d)
$\frac{\partial v_{1}}{\partial t}$ $=-\frac{v_{1}}{R_{1}C_{1}}+\frac{i_{x}}{C_{1}},$ (3e)
$\frac{\partial v_{2}}{\partial t}$ $=-\frac{v_{2}}{R_{2}C_{2}}+\frac{i_{x}}{C_{2}},$ (3f)
where $\Omega=[0,L_{y}]\times[0,L_{z}]$ is the projection of the cell onto the $(y,z)$-plane and $L_{y}$ and $L_{z}$ are the cell width and height, respectively. Eqn. (3a) defines the through-thickness current (as in the current flowing across the electrodes charging the active material particles). Eqn. (3b) defines the distribution of the voltage across the two current collectors. Eqn. (3c) defines the localised temperature dynamics of the cell. Using the circuit model of Figure 1(b) to describe the electrical dynamics at each point ($y,z)$ in the plane, Eqn. (3d) describes the state-of-charge and Eqns. (3e)-(3f) are the relaxation dynamics associated with Figure 1(b)’s RC pairs.
@@ 2.4 Numerical Solution Procedure
It is also noted that graded planar cells were also investigated in Hosseinzadeh, et al. 201841, where a DFN-type model with grading both in-the-plane and through-the-thickness of the electrodes was developed. The two main differences between this paper and Hosseinzadeh, et al. 201841 are: i) an analytic solution for planar graded electrodes to achieve a uniform current distribution is obtained here (see Eqn. (7)), ii) this paper’s model predictions are validated against experimental data from Lin, et al, 202220, iii) it is proposed that the benefits of in-the-plane grading are to reduce electrode degradation (especially plating), whereas voltage and thermal considerations were considered in 41.
@@ 2.5 Comparison to experimental data
The experimental data from Lin, et al, 202220 is used to validate the lithium-ion battery pouch cell model of (3). Briefly, the data is obtained from square-wave-excitation cycling experiments starting at 30% state of charge (SoC), using 20 Ah pouch cells from A123 Systems with LFP positive electrodes and graphite negative electrodes, which were cycled at 4C (at currents ±$\pm 80$ A) every 100 seconds. A thermal imaging camera was then used to capture thermograms of the cell surface and to visualise the cell’s temperature distribution across the plane during the cycling.

### v2

# Graded lithium-ion battery pouch cells to homogenise current distributions and reduce lithium plating

@@ Abstract
Spatial distributions in current, temperature, state-of-charge and degradation across the plane of large format lithium-ion battery pouch cells can significantly impact their performance, especially at high C-rates. In this paper, a method to smooth out these spatial distributions by grading the electrode microstructure in-the-plane is proposed. A mathematical model of a large format pouch cell is developed and validated against both temperature and voltage experimental data. An analytical solution for the optimal graded electrode that achieves a uniform current distribution across the pouch cell is then derived. The model predicts that the graded electrodes could significantly reduce the likelihood of lithium plating in large format pouch cells, with grading increasing the C-rate at which plating occurs from 2.4C to 4.3C. These results indicate the potential of designing spatially varying electrode architectures to homogenise the response of large format pouch cells and improve their high rate performance.
@@ 1 Introduction
Lithium-ion battery electrodes are complex non-linear dynamical systems reacting non-uniformly in space and time, especially when subject to high C-rate currents. Phenomena such as travelling wave fronts in LiFePO4 (LFP) cathodes during charging 1; 2 and localised lithium-plating in graphite anodes 3 are just some examples of the diverse range of non-linear behaviours that have been observed in battery electrodes, and it is these localised behaviours that often govern cell performance. To protect against these spatially localised electrochemical behaviours, the use of graded electrode structures (as in those with controlled spatial variations in electrode microstructure) has been proposed with the idea being that introducing heterogeneity into the electrode microstructure can reduce the heterogeneity in the electrochemical response.
@@
Recently, experimental data have provided support for the use of graded electrodes in practical cells 4. A common theme is the potential of graded electrodes to increase both the capacity at high C-rates 5; 6 and the cycle life of thick electrodes 7; 8, with other benefits including a reduction in electrode tortuosity 9. Data for both standard electrodes, such as LFP cathodes 7; 8, and non-standard electrodes, such as anodes arranged in a layered structure composed of high power Li4Ti5O12 and high capacity SnO2 10, have shown similar benefits. However, care has to be taken when designing graded electrodes, as structuring the electrode incorrectly can deteriorate performance 11; 8. To prevent this, mathematical models have been developed to predict the electrochemical response of graded electrodes so as to quickly and cost-effectively optimise their designs. Several different models have now been developed covering a range of length scales (including micro-structural12; 13 and Doyle-Fuller-Newman-type models14; 2) and performance metrics (such as fast charging capacitance 2 and electrical impedance 15). These results highlight the growing maturity of battery modelling and parameterisation, which in turn, is making it increasingly straightforward to embed modelling pro-actively in the battery design process 16. An example of this battery co-design process are the trapezoidal electrodes7 that were developed using insights 15 on how the distribution of carbon and binder affects the conductivity of the LiFePO4 electrodes.
@@
The focus of most existing studies on graded electrodes for lithium-ion batteries has been on controlling the through-thickness local electrode microstructure 12; 17, including bilayer cathodes2 composed of discrete layers of both     Li[Ni0.6Co0.2Mn0.2]O2 (NMC622) and LFP. Focusing on the through-thickness electrode structure (with experiments typically only involving coin cells 7) is justified by the fact that it is charge mobility through the electrode thickness between the anode and cathode that governs coin cell performance. However, in recent years, there has been a general trend in commercial cells towards larger cell formats (demonstrated, for example, by the introduction of the larger 4680 cylindrical cells 18). The idea being that larger cell formats increase the specific energy density, as the ratio of active material to superfluous materials such as casings and control circuits is increased. Large format cells have their own problems compared to coin cells; in particular, they experience new forms of spatial heterogeneities, which emerge at the macro-scale (as in across their width and length). In particular, heterogenous state-of-charge 25, thermal hot spots 19; 20 and spatially localised distributions of electrode degradation 21; 22; 23, lithium plating 24, and currents in-the-plane of large-format pouch cells 26; 27 have been identified as impacting performance, especially in high rate applications.
@@
The significance of in-the-plane spatial heterogeneities on the performance of large format pouch cells is becoming increasingly apparent, and has driven efforts to understand how these heterogeneities develop using mathematical models. Through these models, predictions for chemical reaction 28 and current distributions along the plane of the current collectors 29 (and how these distributions evolve with changes in the cell’s design, e.g. its current collector thickness 30 and tapering 31) have been developed, with analytical solutions even being proposed 32. The push to understand the planar dynamics of pouch cells has meant that the number of spatial dimensions in these models has increased, with 2D 33 and 3D 34; 35; 36; 37; 20; 38 effects often included, as opposed to the through-thickness models of, for example, the DFN model and single particle models 39; 40. One of the main drawbacks of large format battery models is their complexity, which is a consequence of their multi-dimensional spatial domains. This model complexity leads to several challenges including long simulation times and difficulties in model parameterisation, as discussed in the electro-thermal modelling studies 1920. These challenges demonstrate a need for easily parameterised and computationally efficient pouch cell models able to capture the spatial distributions seen in experimental data. This problem is considered with the model introduced in Section 2.
@@
(a) Representation of electrode/separator sandwich of the pouch cell. A left handed co-ordinate system is used.
@@
Figure 1: Schematic of the 2D model for a lithium-ion battery pouch cell, Figure 1(a) is approximated with Figure 1(b) and Figure 1(c). The model is a combination of Figure 1(b) and Figure 1(c).
The formation of spatial heterogeneities across pouch cells indicates that it may be possible to grade the local electrode microstructure in-the-plane to mitigate their effect, similar to what has been achieved with through-thickness grading 41; 17. The graded electrode problem is illustrated in Figure 1, with Figure 1(a) indicating the planar and through-thickness directions of the pouch cell, Figure 1(b) shows the simplification of the model to a 2D problem with RC circuits distributed across the plane, and Figure 1(c) illustrates the 2D model domain. Figure 1(a) is approximated with Figure 1(b) and Figure 1(c). The model is a combination of Figure 1(b) and Figure 1(c).
@@
    An analytical solution for the distribution of carbon black in the electrode plane that achieves a flat current distribution across the pouch cell is derived.
@@
    Compared to uniform electrodes, it is predicted that graded electrodes can increase the charging rate at which lithium plating occurs from 2.4C to 4.3C.
It is proposed that a primary advantage of these graded pouch cells is to mitigate the impact of lithium plating (see Section 4), as the flatter current distributions reduce the likelihood of negative anode overpotentials. The results highlight the importance of in-the-plane effects when characterising lithium plating in large format cells. By contrast, it is shown that the high thermal/electronic conductivity of the current collectors means that only minimal benefits of grading are observed when considering the thermal and voltage response; Section 3.4 shows that the graded cells only delivered a 1.2% increase in stored capacitance for 4C charging compared to the uniform electrodes, a conclusion in agreement with Hosseinzadeh et al., 201841. These validated modelling results indicate the potential of grading to improve the cycle-life and high-rate performance of large format pouch cells. More broadly, they indicate how improvements in battery design and manufacturing could reduce the performance gap observed between coin cells constructed in laboratories and the large format cells used in commercial high-rate applications 42.
@@ 2 Mathematical model
In this section, a mathematical model of a large format pouch cell is introduced. Specific focus is given to a comparison between the model and experimental data20, with the presented model being simpler in form (specifically, here, circuit dynamics are used to capture the through-thickness electrical response, whereas a simple electrochemical model is used in Lin et al., 202220) whilst still being able to accurately capture the in-plane spatial distributions seen in the experimental data (see Section 2.5). Keeping the model formulation relatively simple while retaining the important physics-based mechanisms, allows analytical solutions to the optimal graded electrode design problem to be obtained (explored in Section 3), which is the focus of this work.
@@ 1. A1:
    The thicknesses of the current collectors are much thinner than their width and height.
@@ 3. A3:
    The multi-layer pouch cell can be modelled as an effective single layer19; 20.
A1 is justified because large format pouch cells are considered (with typical current collector thicknesses being $\approx 10\mu$m and their planar dimensions being $\approx$ 15cm $\times$ 20cm). Defining $i_{x}(y,z,t)$ as the current density through the electrode, and $I_{x,k}(x,y,z,t)$ as the current density flowing in the $x$-direction through current collector $k$, then, under this thin plate assumption, the current density gradient in the $x$-direction through the current collectors can be approximated by
@@
following the ideas of Taheri et al., 201432. A2 is a simplification of the electrode electrical dynamics, but, even so, the resulting model is able to capture the spatial distributions seen in the experimental data (see Section 2.5) while being simple enough to generate an analytical solution to the optimal electrode design problem (see Section 3.1). A3 is justified under the assumption that each layer has a similar composition and so reacts similarly.
@@ 2.2 Model equations
The following equations characterise the proposed mathematical model for large-format pouch cells, with the models domain described in Figure 1(c). Regarding the notation, each model variable in Eqn. (3) is defined in the ($y,z$)-plane of the pouch cell, as illustrated in Figure 1(c), and may also evolve in time $t$. This $(y,z,t)$ dependency of the variables is dropped in the model equations (3) to ease readability. $V(y,z,t)$ is the voltage (defined as the potential of the current collector at the cathode minus that of the anode) at each point in the plane, $T(y,z,t)$ is the cell temperature, $i_{x}(y,z,t)$ is the local through-thickness current density, $R(y,z)$ is the local resistance of the two electrodes and the separator that is constant with time, while SoC$(y,z,t)$ is the local state of charge in the plane. Table A1 summarises the model’s variables, while Table A2 gives its parameters. For the graded cell, the only parameter varying in $(y,z)$ is $R(y,z)$, with the rest being constants. Using the notation $\Delta f=\frac{\partial^{2}f}{\partial y^{2}}+\frac{\partial^{2}f}{\partial z^{2}}$ and defining
@@
$i_{x}$ $=\frac{V-U(\text{SoC,T})-v_{1}-v_{2}}{R\,A_{\scriptscriptstyle\text{cell}}},$ (3a)
$\Delta V$ $=\frac{i_{x}}{\sigma_{\text{eff}}(L_{\text{cn}}+L_{\text{cp}})},\ \ \text{on}\ \ \Omega,\ t>0,$ (3b)
$\rho\frac{\partial T}{\partial t}$ $=\lambda_{\scriptscriptstyle\text{eff}}\;\Delta T+\dfrac{1}{L_{x}\ n_{\text{layers}}}\left[R\;A_{\scriptscriptstyle\text{cell}}\;{i_{x}}^{2}-h_{\text{faces}}\left(T-T_{\text{ref}}\right)\right],\ \ \text{on}\ \ \Omega,\ t>0,$ (3c)
$\frac{\partial\text{SoC}}{\partial t}$ $={\color[rgb]{0,0,0}\frac{A_{\scriptscriptstyle\text{cell}}}{C_{\scriptscriptstyle\text{cell}}}i_{x}},\ \ \text{on}\ \ \Omega,\ t>0,$ (3d)
$\frac{\partial v_{1}}{\partial t}$ $=-\frac{v_{1}}{R_{1}C_{1}}+{\color[rgb]{0,0,0}\frac{A_{\scriptscriptstyle\text{cell}}}{C_{1}}i_{x}},$ (3e)
$\frac{\partial v_{2}}{\partial t}$ $=-\frac{v_{2}}{R_{2}C_{2}}+{\color[rgb]{0,0,0}\frac{A_{\scriptscriptstyle\text{cell}}}{C_{2}}i_{x}},$ (3f)
where $\Omega=[0,\;L_{y}]\times[0,\;L_{z}]$ is the projection of the cell onto the $(y,z)$-plane and $L_{y}$ and $L_{z}$ are the cell width and height, respectively. The physical meaning of Eqns. (3e) - (3f) is as follows. Eqn. (3a) is obtained from the circuit modelsplett2015battery distributed across the plane of the pouch cell in Figure 1(b). Eqn. (3a) defines the through-thickness current density (as in the current density flowing across the electrodes charging the active material particles). Eqn. (3b) is adapted from Taheri et al., 32 to account for the circuit dynamics and defines the distribution of the voltage in the $(y,z)$ plane, given the current density $i_{x}$ through the cell at each point. Eqn. (3c) defines the localised temperature dynamics of the cell and is adapted from Lin et al., 20 with Joule heating defined from Eqn. (3a) instead of electrochemical potentials20. Using the circuit model of Figure 1(b) to describe the electrical dynamics at each point ($y,z)$ in the plane (with the derivations of these dynamics discussed in Plettplett2015battery), Eqn. (3d) describes the state-of-charge and Eqns. (3e) - (3f) are the relaxation dynamics associated with Figure 1(b)’s RC pairs.
@@ 2.4 Numerical Solution Procedure
It is also noted that graded planar cells were also investigated in Hosseinzadeh et al., 201841, where a DFN-type model with grading both in-the-plane and through-the-thickness of the electrodes was developed. The two main differences between this paper and Hosseinzadeh et al., 201841 are: i) an analytical solution for planar graded electrodes to achieve a uniform current distribution is obtained here (see Eqn. (7)), ii) in this paper the model predictions are validated against experimental data from Lin et al., 202220, iii) it is proposed that the benefits of in-the-plane grading are to reduce electrode degradation (especially plating), whereas voltage and thermal considerations were considered in 41.
@@ 2.5 Comparison to experimental data
The experimental data from Lin et al., 202220 is used to validate the lithium-ion battery pouch cell model of (3). Briefly, the data is obtained from square-wave-excitation cycling experiments starting at 30% state of charge (SoC), using 20 Ah pouch cells from A123 Systems with LFP positive electrodes and graphite negative electrodes, which were cycled at 4C (at currents ±$\pm 80$ A) for 100 seconds. A thermal imaging camera was then used to capture thermograms of the cell surface and to visualise the cell’s temperature distribution across the plane during the cycling.